Find the LCM 10xy and 6x^3
Found 2 solutions by arunpaul, JimMarshall:
Answer by arunpaul(104)

10xy = 2*5*x*y
6x^3 = 2*3 *x^3
so LCM will be = 2*3*5*x^3*y = 30 yx^3 (Ans)

Answer by JimMarshall(82)

==>10xy = 2*5*x*y
==>6x^3 = 2*3 *x^3
==> LCM will be = 2*3*5*x^3*y
==> 30 yx^3
Ans